# In-context KV-Cache Eviction for LLMs via Attention-Gate

### v2

@@ 3.3 Attention-Gate
In summary, AG dynamically determines which tokens should be retained, optimizing both KV-Cache usage and attention score computation by utilizing an attention-like structure, Sigmoid activation function, and a threshold-based decision for token eviction.
@@ 3.4 Training Implementation
To train the Attention-Gate (AG) effectively, this section outlines the key components of the training process. For more comprehensive details, please refer to Section 4 and Appendix A.
Eviction Loss. To encourage the eviction of unnecessary tokens, we introduce a dedicated loss function called the Eviction Loss. This loss encourages the model to discard as many tokens as possible, which is defined as:
$\ell_{\text{evict}}=\alpha\cdot\frac{1}{L}\sum_{l=1}^{L}\text{ReLU}\left(\overline{\text{AG}}_{l}-\beta\right),$ (1)
where $L$ is the number of layers, and $\overline{\text{AG}}_{l}$ represents the average output of the AG module in layer $l$. In this formula, $\alpha$ adjusts the intensity of KV-Cache eviction, while $\beta\in[0,\tau]$ ensures that eviction does not become overly aggressive.This loss function works alongside the auto-regressive loss to balance token eviction with maintaining model performance. The loss allows for adaptive eviction across layers and attention-heads. As shown in Figure 1, after training, the model learns to increase token eviction in deeper layers, which is not achievable with training-free methods like H2O (Zhang et al., 2024) or StreamingLLM (Xiao et al., 2024).
Initialization. We initialize the AG parameters using Xavier initialization (Glorot & Bengio, 2010) to provide a stable starting point for learning. Additionally, a small constant $\gamma\geq 0$ is added inside Sigmoid, ensuring that the initial retention probabilities are close to 1. This encourages the model to retain most tokens early in training, allowing it to learn which tokens are important more gradually.
@@ 4 Experiments
This section consists of three main parts. First, we evaluate the performance of AG in two scenarios: continual pre-training (CPT) and supervised fine-tuning (SFT). For CPT, we compare AG with classic methods on performance and eviction rates (Table 2). For SFT, we benchmark AG against vanilla LoRA fine-tuning to highlight its task-specific adaptability (Section 4.2). Second, we provide visualization of selected examples to demonstrate the core characteristics of AG (Section 4.3). Finally, we conduct ablation studies to provide further insights into the effectiveness of AG (Section 4.4).
@@ 4.1 Continual Pre-training
Table 1: Performance comparison of Llama2-7B and various KV-Cache eviction strategies across seven tasks. The table reports accuracy (Acc.) for Llama2-7B and all eviction methods, with Llama2-7B serving as the upper bound for accuracy. Metric %Eviction refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from the KV-Cache. The eviction ratio is fixed at 50% for the baseline methods, including a local strategy (retaining only recent tokens), StreamingLLM, and H2O. In contrast, our method, which includes continual pre-training, achieves higher average accuracy while maintaining a higher average %Eviction.
@@
%Eviction 54.29 51.03 51.05 46.70 40.02 57.75 52.16 51.87
Table 2: Performance of Llama2-7B with LoRA fine-tuning and our method on six downstream tasks. In addition to the LoRA fine-tuning targets, our method makes the AG modules learnable. Two settings for $\alpha$ (0.5 and 1) are tested. Our method maintains comparable or better accuracy while achieving a higher eviction ratio, demonstrating its task-specific adaptability in managing token eviction without significant accuracy loss.
Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
Fine-tuned Llama2-7B Acc. 82.92 60.34 64.98 92.00 88.10 78.80 77.86
Ours $(\alpha=1)$ Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
Ours $(\alpha=0.5)$ Acc. 81.50 57.63 74.01 95.00 87.00 79.20 79.06
%Eviction 64.96 36.45 62.80 34.77 67.31 66.65 55.49
4.1.1 Setup
Models and datasets. We use Llama2-7B (Touvron et al., 2023) as our primary base model due to its strong performance and popularity within the AI community. Additionally, we validate the feasibility of our approach on Mistral-7B (Jiang et al., 2023), with results provided in Table 4. For continual pre-training, we select a subset of the Redpajama dataset (Computer, 2023) containing approximately 5,000 samples  [^1] We sampled 4,997 samples proportionally from each subset of RedPajama.. To assess the effectiveness of our method, we evaluate it on seven widely recognized benchmarks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and HellaSwag (Zellers et al., 2019). All evaluations are conducted in a zero-shot setting, with accuracy assessed using OpenCompass (Contributors, 2023).
Training details. During continual pre-training (CPT), only the AG modules are learnable, while the rest of the model’s parameters are frozen. For the Llama2-7B model, the hyperparameters from Section 3.4 are used: $\tau=0.5$, $\gamma=2$, $\alpha=5$, and $\beta=0.4$. The model was trained for one epoch.
Metrics and Baselines. The metrics considered include accuracy and the mean eviction ratio for all KV-Cache. The accuracy of vanilla Llama2-7B on these datasets is used as the upper bound, with the goal of minimizing accuracy loss while maximizing the mean eviction ratio. For KV-Cache eviction strategies, we select a local method (retaining only recent tokens) and StreamingLLM (Xiao et al., 2024) as representatives of static strategies, while H2O (Zhang et al., 2024) serves as a representative of methods based accumulative attention scores. For all these methods, eviction ratio is set to 50%.
4.1.2 Results
The results in Table 1 demonstrate the effectiveness of our method in balancing accuracy and KV-Cache eviction. Our method consistently outperforms the baseline strategies in terms of accuracy across most tasks while achieving a higher mean eviction ratio. This indicates that our method is able to evict more tokens on average without significantly compromising accuracy.
In terms of accuracy, our method achieves competitive performance, closely matching the performance of vanilla Llama2-7B in most cases. For instance, on PIQA, COPA, and BoolQ, our method performs comparably to Llama2-7B, demonstrating minimal accuracy degradation. Compared to baselines, our method shows superior performance, particularly in tasks like BoolQ and PIQA, where both accuracy and eviction ratio surpass the baselines. This confirms the advantage of learnable mechanism for more efficient token retention strategies without sacrificing model performance.
Moreover, it is worth highlighting that our method achieves these results with minimal computational overhead, as mentioned in Section 4.1.1. The continual pre-training was conducted on only 5,000 samples and trained for just one epoch, demonstrating the lightweight nature of our approach. This efficiency can be attributed to the fact that our method does not need to learn new knowledge from scratch but rather focuses on learning effective token retention strategies, leveraging the existing capabilities of the pre-trained model.
4.2 Supervised Fine-tuning
Figure 3: This visualization highlights attention patterns in Llama2-7B after fine-tuning on the BoolQ dataset, using a selected sample to showcase multiple heads within both MHA and AG of different layers. In part (i), we visualize attention scores from several MHA heads across different layers before eviction. 1. MHA heads exhibit diverse attention patterns, especially in the first two layers, where attention is highly heterogeneous across heads. 2. While the first two layers show dense attention patterns, the subsequent layers become progressively sparser. 3. Bright-yellow vertical lines consistently appear at the same position across nearly all heads, especially beyond the first two layers, indicating critical tokens for inference. This aligns with the Heavy Hitters in H2O (Zhang et al., 2024), where a small portion of tokens significantly contributes to the attention scores. Our method’s effectiveness is demonstrated by the persistence of these bright lines in deeper layers, showing that crucial tokens are retained rather than evicted, ensuring their importance is preserved across the network. In part (ii), we visualize the attention scores from the attention-like structure within the AG mechanism. As the model layers deepen, the attention pattern shifts from high-resolution to lower-resolution, indicating that AG increasingly focuses on distilling in-context information. Thus, deeper AG layers no longer need such high resolution to capture global information, as the model has already refined and summarized it. This suggests potential optimizations, such as reducing the number of heads or dimensionality in deeper AG layers, to further improve efficiency.
4.2.1 Setup
Model and tasks. We utilize Llama2-7B (Touvron et al., 2023) as base model. To evaluate our approach, we select six widely recognized downstream tasks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and OpenBookQA (Mihaylov et al., 2018). For each task, we fine-tune the model using the respective training set and evaluate its performance on the corresponding test set.
Baselines and Implementation Details. For the baseline, we apply LoRA (Hu et al., 2021) to fine-tune the models on each task, targeting the $(W_{q},W_{k},W_{v},W_{o})$ weights in the self-attention modules. In our method, in addition to applying LoRA to these parameters, the AG modules are also set to be learnable. As mentioned in Section 3.4, the following hyperparameters are used: $\tau=0.3$, $\gamma=0$, $\alpha=1$ or $0.5$, and $\beta=0.28$. We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5, and train for two epochs on each dataset.
4.2.2 Results
As shown in Table 2, our method demonstrates a strong balance between accuracy and KV-Cache eviction across the six downstream tasks.
With $\alpha=1$, our method maintains competitive accuracy compared to the fine-tuned Llama2-7B baseline, with minimal drops (e.g., 82.15% vs. 82.92% on PIQA), while achieving a high mean eviction ratio of 60.00%. With $\alpha=0.5$, the eviction rate is reduced to 55.49%. But in some tasks like RTE and COPA, our method even surpasses the baseline accuracy. The average accuracy also exceeds the baseline (79.06% vs. 77.86%), suggesting that effectively evicting redundant tokens allows the model to focus on relevant information and improves performance.
Additionally, under the same hyperparameter settings, the performance varies across tasks. For instance, ARC-C is more challenging to evict compared to OpenBookQA, leading to a larger accuracy drop post-eviction. This highlights the importance of task-specific KV-Cache eviction policies.
4.3 Visualization
Table 3: Ablation study on various settings of AG, reporting accuracy (Acc.) and KV-Cache eviction ratio (%Eviction) under different configurations, with $\alpha=1$ in all settings. The results for (1) correspond to the setup described in Section 4.2, where the number of AG heads is 4, and the head dimension is 128. In (2-1) and (2-2), we explore the impact of the number of AG heads, with (2-1) using 2 heads and (2-2) using 1 head. The comparison between (1), (2-1), and (2-2) shows that reducing the number of heads leads to a drop in both accuracy and eviction ratio, indicating that the capacity of AG is closely tied to the number of heads. For (3-1) and (3-2), we assess the effect of reducing head dimensions for AG heads, where (3-1) has half the dimension size of (1) and (3-2) has 1/4. Comparing (1), (3-1), and (3-2) reveals that smaller dimensions reduce the eviction capability and accuracy, highlighting the importance of maintaining sufficient head dimensionality. Settings (4-1) and (4-2) examine using the hidden states and the AG module from the previous layer to inform the current layer’s eviction strategy. In (4-1), the first layer does not evict, and from the second layer onward, the eviction is determined by the previous layer. In (4-2), the first and second layers do not evict, following the observation from Section 4.3 that the first two layers are denser than subsequent layers. Both (4-1) and (4-2) show a slight decline in accuracy and eviction ratio compared to (1), but this approach introduces parallelism, offering a potential avenue for future optimizations. The setting (5) replaces the attention-like structure in AG with a simple linear layer to determine the eviction strategy. The comparison with (1) shows that linear layers almost cannot evict tokens effectively, reinforcing the necessity of leveraging global in-context information for successful eviction, as discussed in Section 3.3.
Metric PIQA ARC-C RTE COPA BoolQ OpenBookQA Avg.
(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Eviction 66.16 48.31 65.47 45.40 67.46 67.17 60.00
(2-1) Acc. 81.88 57.63 65.70 91.00 87.52 77.40 76.86
%Eviction 63.92 36.38 62.73 24.38 65.22 63.57 52.70
(2-2) Acc. 82.15 53.90 62.45 89.00 87.31 77.40 75.37
%Eviction 58.97 31.47 59.77 20.32 63.02 59.17 48.79
(3-1) Acc. 81.45 53.36 58.84 88.00 86.73 78.40 74.46
%Eviction 61.75 33.55 61.34 19.24 64.59 59.59 50.01
(3-2) Acc. 83.03 53.90 59.93 89.00 87.16 76.40 74.90
%Eviction 58.68 24.23 32.23 12.28 59.40 55.54 40.39
(4-1) Acc. 81.66 55.25 66.06 88.00 86.85 78.00 75.97
%Eviction 49.52 36.92 46.85 28.74 56.02 60.32 46.40
(4-2) Acc. 82.75 55.93 79.06 82.00 86.33 78.40 77.41
%Eviction 53.31 44.38 51.20 47.95 61.98 61.73 53.43
(5) Acc. 82.54 54.58 57.40 81.00 87.71 74.80 73.01
%Eviction 1.06 0.46 0.81 0.26 1.38 1.16 0.86
In this section, we present visualizations to highlight key characteristics of our AG mechanism. After fine-tuning on specific tasks, we visualize the model’s MHA and AG weights for selected samples, as shown in Figure 3. Additional visualizations can be found in Section A.2.
4.4 Ablation
In this ablation study, we explore the effects of various configurations on the AG mechanism, focusing on the number of AG heads, AG head dimensions, and eviction strategies. For details of each setting and the corresponding performance, refer to Table 3. Key findings include the impact of reducing the number of heads, as shown in (2-1) and (2-2), and the reduction of head dimensions in (3-1) and (3-2), both of which result in lower accuracy and eviction ratios. In (4-1) and (4-2), we examine eviction strategies where the current layer’s eviction is based on the previous layer’s hidden states and AG module, which introduces parallelism but slightly reduces performance. Finally, replacing the AG with a linear layer in (5) demonstrates the necessity of using an attention-like structure for effective token eviction. For more ablation, please refer to Table 5.
5 Conclusion
In conclusion, the proposed Attention-Gate mechanism offers a flexible and adaptive solution to KV-Cache eviction in large language models. By dynamically identifying and discarding less important tokens, Attention-Gate addresses the limitations of static and attention-score-based strategies, providing efficient context-aware eviction. This mechanism integrates seamlessly with pre-trained models and can be easily tuned, making it a practical and effective method for enhancing both performance and memory efficiency in various tasks.
Appendix A Appendix
A.1 Results of Continual Pre-training on Mistral

### v3

@@ 3.5 Training Implementation
To train the Attention-Gate (AG) effectively, this section outlines the key components of the training process. For more details, please refer to Section 4 and Appendix A.
Eviction Loss. To encourage the eviction of unnecessary tokens, we introduce a dedicated loss function called the Eviction Loss. The loss function encourages the model to maintain the eviction ratio close to a target value $\beta$, which is defined as:
$\ell_{\text{evict}}=\alpha\cdot\left|\overline{\text{AG}}-\beta\right|,$ (6)
where $\overline{\text{AG}}$ represents the average output of all AG modules. In this formula, $\alpha$ adjusts the intensity of KV-Cache eviction, while $\beta\in[0,\tau]$ ensures that eviction does not become overly aggressive. Eviction Loss allows for adaptive eviction across layers and attention-heads. This loss function works alongside the auto-regressive loss to balance token eviction with maintaining model performance.
Initialization. We initialize the AG parameters using Xavier initialization (Glorot & Bengio, 2010) to provide a stable starting point for learning. Additionally, a small constant $\gamma\geq 0$ can optionally be added inside Sigmoid in Equation 4, ensuring that the initial retention probabilities are close to 1. This encourages the model to retain most tokens early in training.
@@
Table 1: Performance comparison of Llama2-7B and various KV-Cache eviction strategies after continual pre-training. For baselines, ($W_{q},W_{k},W_{v},W_{o}$) are made trainable, while in our method, the AG module is also trainable. Higher values indicate better performance for all metrics. Acc. refers to accuracy. %Evict. refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from KV-Cache. The eviction ratio is fixed at 50% for the baseline methods. In contrast, our method achieves better performance (average accuracy and score) while maintaining a higher average %Evict..
Metric PIQA ARC-C ARC-E RTE COPA BoolQ HellaSwag MMLU Avg. Metric LongBench
Llama2-7B-cpt Acc. 72.69 32.88 50.62 50.54 57.00 64.77 42.19 26.64 49.67 Score 23.42
StreamingLLM Acc. 72.42 31.53 49.74 50.90 54.00 61.31 37.75 26.66 48.04 Score 4.61
H2O Acc. 72.20 30.85 49.38 51.99 55.00 62.42 41.45 26.45 48.72 Score 4.85
Ours Acc. 76.33 32.20 48.32 50.18 59.00 60.46 64.23 28.54 52.41 Score 13.71
%Evict. 43.12 46.54 45.15 48.60 55.37 50.16 61.10 70.36 52.55 %Evict. 68.55
3.6 Complexity Analysis
AG module. For input $X\in\mathbb{R}^{n\times d}$, the FLOPs for AG are:
$\operatornamewithlimits{\text{FLOPs}}_{\text{AG}}=\mathcal{O}(n^{2}d_{k}^{\prime}h^{\prime})\ .$
MHA module. Without AG, original MHA’s FLOPs are:
$\operatornamewithlimits{\text{FLOPs}}_{\text{original MHA}}=\mathcal{O}(n^{2}d_{k}h)\ .$ (7)
After AG processing, $t\%$ of the KV-Cache tokens are discarded, leaving $(1-t\%)$ for attention computation:
$\operatornamewithlimits{\text{FLOPs}}_{\text{MHA after AG}}=\mathcal{O}\big{(}(1-t\%)n^{2}d_{k}h\big{)}\ .$
Combined AG & MHA, the total FLOPs are:
$\operatornamewithlimits{\text{FLOPs}}_{\text{AG \& MHA}}=\mathcal{O}\big{(}n^{2}\max\big{(}d_{k}^{\prime}h^{\prime},(1-t\%)d_{k}h\big{)}\big{)}\ .$ (8)
Efficiency. The reduction in FLOPs depends on three factors: (i) Reduction in token count ($t\%$): Higher values of $t\%$ result in a larger reduction in the quadratic term of the original MHA. (ii) Head configuration ($h^{\prime}<h$): The AG module must have significantly fewer heads ($h^{\prime}$) compared to the original MHA ($h$) to ensure its overhead is small. (iii) Head dimension ratio ($d_{k}^{\prime}<d_{k}$): A smaller head dimension ($d_{k}^{\prime}$) for AG further reduces its contribution to total FLOPs.
@@ 4 Experiments
This section consists of three main parts. First, we evaluate the performance of AG in two scenarios: continual pre-training (CPT) and supervised fine-tuning (SFT) (Section 4.1 & 4.2). Second, we provide visualization of selected examples to demonstrate the core characteristics of AG (Section 4.3). Finally, we conduct ablation studies to provide further insights into the effectiveness of AG (Section 4.4). Additional results are provided in Section A.1.
@@ 4.1 Continual Pre-training
4.1.1 Setup
Models & datasets. We use Llama2-7B (Touvron et al., 2023) as our primary base model. Additionally, we validate the feasibility of our approach on Mistral-7B (Jiang et al., 2023), with results provided in Table 5. For continual pre-training, we select a subset of the RedPajama dataset (Computer, 2023), comprising approximately 5,000 samples  [^1] Specifically, we sampled 4,997 samples proportionally from each subset of the RedPajama dataset., to serve as the training set. To assess the effectiveness of our method, we evaluate it on widely recognized benchmarks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021) and LongBench (Bai et al., 2023). All evaluations are conducted in a zero-shot setting, with performance assessed using OpenCompass (Contributors, 2023).
Training details & baselines. For the Llama2-7B model, the following hyperparameters from Section 3.5 were used: $\tau=0.5$, $\gamma=2$, $\alpha=5$, and $\beta=0.4$. The model was trained for a single epoch. The evaluation metrics include performance (accuracy and score) and the mean eviction ratio for all KV-Cache. For KV-Cache eviction strategies, we use StreamingLLM (Xiao et al., 2024) as a representative of static strategies, while H2O (Zhang et al., 2024) represents methods based on accumulative attention scores  [^2] Adams et al. (2024) suggest that H2O is an optimal solution in many scenarios. Due to its strong performance and the difficulty of reproducing other training-free methods, no additional training-free baselines were included.. For baseline methods, eviction ratio is fixed at 50%, and ($W_{q},W_{k},W_{v},W_{o}$) are made trainable using LoRA (Hu et al., 2021). In our method, AG is also trainable  [^3] We tested a version where only AG is trainable, with other parameters frozen. Results are shown in Table 4..
4.1.2 Results
The results in Table 1 demonstrate the effectiveness of our method in balancing performance and KV-Cache eviction. Our method consistently outperforms the baseline strategies in terms of performance across most tasks while achieving a higher mean eviction ratio.
Moreover, it is worth highlighting that our method achieves these results with minimal computational overhead, as detailed in Section 4.1.1. The continual pre-training was conducted on only 5,000 samples and trained for just one epoch, demonstrating the lightweight nature of our approach. This efficiency can be attributed to the fact that our method does not need to learn new knowledge from scratch but rather focuses on learning effective token retention strategies, leveraging the existing capabilities of the pre-trained model.
Table 2: Performance of Llama2-7B with LoRA fine-tuning and our method on six downstream tasks. In addition to the LoRA fine-tuning targets, our method makes the AG modules learnable. Two settings for $\alpha$ (0.5 and 1) are tested. Our method maintains comparable or better accuracy while achieving a higher eviction ratio, demonstrating its task-specific adaptability in managing token eviction.
Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
Fine-tuned Acc. 82.92 60.34 64.98 92.00 88.10 78.80 77.86
Llama2-7B
Ours Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
$(\alpha=1)$Evict. 66.16 48.31 65.47 45.40 67.46 67.17 60.00
Ours Acc. 81.50 57.63 74.01 95.00 87.00 79.20 79.06
$(\alpha=0.5)$Evict. 64.96 36.45 62.80 34.77 67.31 66.65 55.49
4.2 Supervised Fine-tuning
4.2.1 Setup
Model & tasks. We utilize Llama2-7B (Touvron et al., 2023) as base model. To evaluate our approach, we select six widely recognized downstream tasks: PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), RTE (Bar-Haim et al., 2006), COPA (Roemmele et al., 2011), BoolQ (Clark et al., 2019), and OpenBookQA (Mihaylov et al., 2018). For each task, we fine-tune model using the respective training set and evaluate its performance on the corresponding test set.
Implementation details. The selection of trainable parameters follows Section 4.1.1. The hyperparameters are $\tau=0.3$, $\gamma=0$, $\alpha=1$ or $0.5$, and $\beta=0.28$. Training is performed using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 5e-5 for 2 epochs per dataset.
4.2.2 Results
As shown in Table 2, our method achieves a strong balance between accuracy and KV-Cache eviction.
With $\alpha=1$, it maintains competitive accuracy compared to the fine-tuned Llama2-7B baseline while achieving a high mean eviction ratio of 60.00%. With $\alpha=0.5$, the eviction ratio decreases to 55.49%, but the average accuracy improves . In tasks like RTE and COPA, it even surpasses the baseline. This suggests that effective token eviction helps the model focus on relevant information.
Additionally, performance varies across tasks under the same settings. For instance, ARC-C is more challenging to evict compared to OpenBookQA, leading to a larger accuracy drop post-eviction. This highlights the importance of task-specific KV-Cache eviction policies.
Figure 3: Comparison of peak memory usage and prefilling time between the LLaMA2-7B model (without AG) and the proposed implementation (with AG and $\sim$50eviction) across varying prompt lengths. The results show significant improvements in memory efficiency with AG, especially as prompt length increases. Prefilling time is not the primary focus, and the current implementation (marked with * in the legend) relies on a suboptimal for-loop over attention heads. Even so, the method maintains stable prefilling time and shows a clear reduction trend with longer prompts.
Figure 4: This visualization highlights attention patterns in Llama2-7B after fine-tuning on the BoolQ dataset, showcasing multiple heads within both MHA and AG across different layers using a selected sample. In part (i), we visualize attention scores from several MHA heads across layers before eviction: 1. MHA heads display diverse attention patterns, especially in the first two layers, where heterogeneity is prominent. 2. Attention patterns become progressively sparser in deeper layers, transitioning from dense in early layers. 3. Bright-yellow vertical lines, indicating critical tokens for inference, appear consistently across heads in deeper layers. These align with the Heavy Hitters in H2O (Zhang et al., 2024), highlighting tokens that significantly contribute to attention scores. Our method ensures these critical tokens are preserved in deeper layers, maintaining their importance across the network. In part (ii), we visualize the attention-like scores from the AG mechanism: As layers deepen, AG shifts from high-resolution to lower-resolution attention, focusing on distilling in-context information. Deeper AG layers no longer require high resolution for capturing global context, as earlier layers have already refined it. This suggests potential optimizations, such as reducing the number of heads or dimensions in deeper AG layers, to enhance efficiency.
4.2.3 Inference Efficiency
We evaluated the inference efficiency of the LLaMA2-7B model with and without the AG module. As shown in Figure 3, our method achieves notable improvements in memory efficiency, especially with longer prompts. Although prefilling time is influenced by the current implementation, it remains stable and demonstrates scalability as prompt lengths increase. We believe there is room for acceleration using kernel fusion techniques and leave it for future work.
4.3 Visualization
Table 3: Ablation study on various settings of AG, reporting accuracy (Acc.) and KV-Cache eviction ratio (%Evict.) under different configurations, with $\alpha=1$ in all settings. The results for (1) correspond to the setup described in Section 4.2, where the number of AG heads is 4, and the head dimension is 128. In (2-1) and (2-2), we explore the impact of the number of AG heads, with (2-1) using 2 heads and (2-2) using 1 head. The comparison between (1), (2-1), and (2-2) shows that reducing the number of heads leads to a drop in both accuracy and eviction ratio, indicating that the capacity of AG is closely tied to the number of heads. For (3-1) and (3-2), we assess the effect of reducing head dimensions for AG heads, where (3-1) has half the dimension size of (1) and (3-2) has 1/4. Comparing (1), (3-1), and (3-2) reveals that smaller dimensions reduce the eviction capability and accuracy, highlighting the importance of maintaining sufficient head dimensionality. Settings (4-1) and (4-2) explore using the previous layer’s hidden states and AG module to guide the current layer’s eviction strategy. In (4-1), eviction begins from the second layer onward, while in (4-2), eviction starts from the third layer, reflecting the denser attention patterns in the first two layers (Section 4.3). Both settings show a slight decline in accuracy and eviction ratio compared to (1) but introduce parallelism, suggesting potential for future optimization. The setting (5) replaces MHA in AG with a simple linear layer to determine the eviction strategy. The comparison with (1) shows that linear layers almost cannot evict tokens effectively, reinforcing the necessity of leveraging global in-context information for successful eviction, as discussed in Section 3.4.
Metric PIQA ARC-C RTE COPA BoolQ OBQA Avg.
(1) Acc. 82.15 59.66 64.26 93.00 86.82 78.80 77.45
%Evict. 66.16 48.31 65.47 45.40 67.46 67.17 60.00
(2-1) Acc. 81.88 57.63 65.70 91.00 87.52 77.40 76.86
%Evict. 63.92 36.38 62.73 24.38 65.22 63.57 52.70
(2-2) Acc. 82.15 53.90 62.45 89.00 87.31 77.40 75.37
%Evict. 58.97 31.47 59.77 20.32 63.02 59.17 48.79
(3-1) Acc. 81.45 53.36 58.84 88.00 86.73 78.40 74.46
%Evict. 61.75 33.55 61.34 19.24 64.59 59.59 50.01
(3-2) Acc. 83.03 53.90 59.93 89.00 87.16 76.40 74.90
%Evict. 58.68 24.23 32.23 12.28 59.40 55.54 40.39
(4-1) Acc. 81.66 55.25 66.06 88.00 86.85 78.00 75.97
%Evict. 49.52 36.92 46.85 28.74 56.02 60.32 46.40
(4-2) Acc. 82.75 55.93 79.06 82.00 86.33 78.40 77.41
%Evict. 53.31 44.38 51.20 47.95 61.98 61.73 53.43
(5) Acc. 82.54 54.58 57.40 81.00 87.71 74.80 73.01
%Evict. 1.06 0.46 0.81 0.26 1.38 1.16 0.86
In this section, we present visualizations to highlight key characteristics of our AG mechanism. After fine-tuning on specific tasks, we visualize the model’s MHA and AG weights for selected samples, as shown in Figure 4. The complete version can be found in Section A.3.
4.4 Ablation
In the ablation study, we investigate the effects of various configurations on the AG mechanism, including the number of AG heads, head dimensions, and eviction strategies. Detailed settings and performance results are provided in Table 3. Key findings include the impact of reducing the number of heads (as seen in (2-1) and (2-2)) and head dimensions (in (3-1) and (3-2)), both of which lead to lower accuracy and eviction ratios. In (4-1) and (4-2), we evaluate eviction strategies where the current layer’s eviction is based on the previous layer’s hidden states and AG module, introducing parallelism but with a slight performance trade-off. Replacing MHA with a linear layer in setting (5) highlights the importance of an attention-like structure for effective token eviction. Additional results are provided in Table 6.
5 Conclusion
In conclusion, the proposed Attention-Gate mechanism offers a flexible and adaptive solution to KV-Cache eviction in large language models. By dynamically identifying and discarding less important tokens in a data-driven manner, Attention-Gate addresses the limitations of static and attention-score-based strategies, providing efficient context-aware eviction. This mechanism integrates seamlessly with pre-trained models and can be easily tuned, making it a practical and effective method for enhancing both performance and memory efficiency in various tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experiments
A.1 Additional Results for Continual Pre-training
In this section, we perform continual pre-training on Llama2-7B using the same training data and hyperparameter settings described in Section 4.1.1. The baselines are training-free, while in our method, only the AG module is trainable. The results are shown in Table 4.
@@ A.1 Additional Results for Continual Pre-training
Table 4: Performance comparison of Llama2-7B and various KV-Cache eviction strategies across seven tasks. Our approach trains only the AG module during continual pre-training, keeping other components frozen. The table reports accuracy (Acc.) for Llama2-7B and all eviction methods, with Llama2-7B serving as the upper bound for accuracy. Metric %Evict. refers to the mean KV-Cache eviction ratio, representing the percentage of tokens evicted from the KV-Cache. The eviction ratio is fixed at 50% for the baseline methods, including a local strategy (retaining only recent tokens), StreamingLLM, and H2O. In contrast, our method achieves higher average accuracy while maintaining a higher average %Evict..
@@
%Evict. 54.29 51.03 51.05 46.70 40.02 57.75 52.16 51.87
A.2 Results of Continual Pre-training on Mistral